# Toxicity Red-Teaming: Benchmarking LLM Safety in Singapore’s Low-Resource Languages

### v1

@@ 1 Introduction
The rapid proliferation of large language models (LLMs) has introduced unprecedented capabilities in natural language processing, enabling applications ranging from conversational Artificial Intelligence (AI) to multilingual content generation Dwivedi et al. (2023); Gottlieb et al. (2023); Deng et al. (2023). However, alongside these advancements, concerns regarding AI safety, ethical deployment, and content moderation have grown, particularly in managing harmful or toxic content Wang and Kim (2023); Anjum and Katarya (2024). As LLMs are increasingly integrated into real-world applications, ensuring their responsible operation across diverse linguistic and cultural settings is paramount.
Red-teaming approaches have emerged as a prominent method for systematically probing LLM vulnerabilities Perez et al. (2022); Ganguli et al. (2022). Techniques such as adversarial input crafting Papernot et al. (2016), scenario-based evaluations Carroll (1997), and boundary condition testing Papanastasiou et al. (1992) have proven effective for detecting biases, toxic content, and other risks. However, these efforts have been largely concentrated on high-resource languages, such as English, thus leaving critical gaps in low-resource multilingual settings. Limited data availability, combined with unique linguistic variability and cultural nuances, exposes models to new vulnerabilities, increasing the risk of biased and toxic content generation in low-resource multilingual contexts. Such vulnerabilities can be exploited to perpetuate social biases and exacerbate social divides within marginalized communities.
Singapore presents a uniquely relevant testbed for evaluating LLM safety in low-resource, multilingual contexts. The country’s linguistic landscape encompasses English, Singlish (a creole-like vernacular), Chinese, Malay, and Tamil, with frequent code-switching, culturally specific expressions, and sociolectal variations. These local linguistic factors present new challenges for AI safety evaluation that go beyond traditional toxicity detection frameworks, which were primarily designed for English. While recent efforts, such as SeaLLM Zhang et al. (2024) and SEA-LION Singapore (2024), have developed LLMs tailored to Southeast-Asian languages, the effectiveness of their safety mechanisms in minority languages remain underexplored. Existing benchmarks, such as SGHateCheck Ng et al. (2024), provide valuable functional tests for hate speech detection, but do not systematically evaluate LLM vulnerabilities under adversarial attacks. Given the increasing adoption of AI-driven communication tools in Singapore and similar multilingual societies, developing a rigorous red-teaming evaluation framework is essential for advancing inclusive and contextually aware AI safety research.
@@
Our work makes the following key contributions: (i) We present SGToxicGuard, the first multilingual dataset explicitly designed for red-teaming LLMs in low-resource linguistic environments. (ii) We propose a three-task evaluation framework —- Conversational Safety, Toxic QA, and Content Composition —- that systematically assesses LLMs’ ability to handle Singapore-specific toxic content in real-world applications. (iii) Our study offers the first large-scale benchmarking of multilingual LLMs across Singapore’s low-resource languages, highlighting strengths and limitations in existing AI safety implementations.
By demonstrating the vulnerabilities of LLMs for multilingual toxicity detection, we lay the groundwork for future improvements in AI safety mechanisms, particularly in linguistically diverse societies. Our findings contribute to ongoing discussions on AI ethics, content moderation policies, and robustness in multilingual AI deployments.
@@ 2.1 Red-Teaming for AI Safety
Singapore’s linguistic landscape, which includes English, Singlish, Mandarin, Malay, and Tamil, introduces unique challenges such as code switching and cultural specificity, which existing frameworks often overlook. We address this gap by proposing a Singapore-focused red-teaming methodology to benchmark LLM safety across multilingual, real-world scenarios, contributing toward more inclusive and resilient AI moderation systems.
@@ 2.2 Multilingual Toxicity Detection
Recent advances in multilingual embeddings and LLMs, such as mBERT Pires (2019), XLM-RoBERTa Conneau (2019), GPT-3 Achiam et al. (2023), LLaMA Dubey et al. (2024), and Mistral Jiang et al. (2023), have improved cross-lingual transfer, but also amplified concerns about bias Tedeschi et al. (2024). We extend this line of research by evaluating LLM toxicity detection using SGHateCheck Ng et al. (2024), a curated dataset offering parallel toxic content across Singlish, Chinese, Malay, and Tamil. Unlike prior Southeast Asian datasets Maity et al. (2023); Lu et al. (2023); Chakravarthi et al. (2021), SGHateCheck enables structured multilingual comparisons, revealing disparities in LLM safety and informing the development of culturally sensitive AI moderation frameworks. Examples of localized hateful content containing local slurs and expressions across Singlish, Chinese, Malay, and Tamil can be found in Appendix A.6.
Beyond SGHateCheck, a growing body of work has advanced hate speech detection across languages and modalities. For example, datasets and benchmarks such as RabakBench Chua et al. (2025) highlight the importance of localized evaluations for low-resource settings. Recent models like LionGuard 2 Tan et al. (2025) and AngryBERT Awal et al. (2021) aim to build lightweight, emotion- and target-aware hate speech detectors, while augmentation approaches such as HateGAN Cao and Lee (2020) and deep representation methods like DeepHate Cao et al. (2020) improve robustness under data scarcity. Other studies emphasize adversarial resilience, e.g., ToxiCloakCN for cloaked Chinese offensive text Xiao et al. (2024), as well as model-agnostic meta-learning approaches for multilingual adaptation Awal et al. (2023). Complementary research explores explainability and moderation workflows, including retrieval-based methods for covert toxicity detection Lee et al. (2024). Together with surveys on multimodality and large models in content moderation Hee et al. (2024) and analyses of Singapore-focused social media hate speech Haber et al. (2023), these efforts demonstrate the breadth of strategies for building culturally grounded, robust hate speech moderation systems.
@@ 3 Framework
3.1 Toxic Conversation Task
Toxic Conversation Task evaluates LLM’s ability to generate non-toxic responses in a single-turn dialogue setting involving hateful input. This task examines the model’s behavior across various social groups and low-resource languages. It simulates a dialogue scenario between two participants, A and B, where participant A provides a hateful statement, and participant B responds. Hateful statements from participant A are sourced from the SGHateCheck dataset Ng et al. (2024), while the LLM generates participant B’s response, which is expected to avoid toxicity by adopting constructive and respectful tones, such as refusal or dissuasion.
@@ 3.3 Toxic Tweet Composition Task
This task was evaluated under both zero-shot and two-shot in-context learning settings. In the two-shot setting, demonstration examples were provided, each showing a transformation of a hateful statement into a tweet that subtly conveys the same hateful message. These examples, while crafted to simulate malicious intent, are used to assess the LLM’s vulnerability to generating subtle hateful content. Additionally, they help evaluate whether exposure to such examples could trigger a “toxicity jailbreak”, where the model deviates from its intended ethical safeguards. Examples of tweet composition and demonstration are listed in Appendix A.2 and A.3.
@@ 5.1 Toxic Conversation Task
We can see in Table 2 that, in the zero-shot setting, most LLMs demonstrate robust safety mechanisms, producing relatively few hateful responses across different languages and the models exhibit lowest hateful response rate in English. However, notable exceptions include Mistral Jiang et al. (2023) and SeaLLM, which exhibit vulnerabilities in specific languages. Mistral generates a sizable proportion of hateful responses in Malay (9.74%) and Tamil (7.86%), underscoring potential weaknesses in handling low-resource languages. Surprisingly, the specialized multilingual LLM, SeaLLM, performs worse, with higher rates of hateful responses in Singlish (8.82%), Chinese (7.46%), and Tamil (14.64%). These results indicate that models fine-tuned for Southeast Asian languages face considerable challenges in mitigating toxicity in low-resource contexts.
@@
Introducing two hateful dialogue demonstrations in the two-shot setting results in a notable increase in the generation of hateful responses across all LLMs and languages. Qwen Team (2024) demonstrates the highest resilience to in-context learning, maintaining the lowest overall rate of hateful responses. However, even Qwen exhibits a sizable increase in Tamil (22.72%). In contrast, Mistral Jiang et al. (2023) and Llama-3.1 Dubey et al. (2024) show moderate increase in hateful response rates across languages, ranging from approximately 2% to 14%. SeaLLM, despite higher rates in the zero-shot setting, demonstrates relatively lower rates in Malay (2.66%) and Chinese (11.01%) under two-shot conditions. SEA-LION Singapore (2024), however, exhibits alarming susceptibility to toxicity induction, with hateful response rates increasing to 61.52% in Chinese and 58.42% in Tamil. These results suggest that in-context learning substantially amplifies the risk of toxicity jailbreaks in certain models.
@@
Our result reveals a major increase in toxic content generation by LLMs when interacting with users of low-resource languages, particularly under adversarial prompting conditions. In zero-shot evaluations, we observed elevated toxicity levels in Malay, Tamil, and even Chinese, suggesting inherent limitations in the safety alignment of multilingual LLMs for these languages. This trend was further amplified in two-shot scenarios, with substantial increase in toxicity rates for Tamil in SEA-LION Singapore (2024) and Qwen Team (2024). Such differences suggest that existing safeguards may be more thoroughly tuned for English data. However, the consistently low toxicity rates observed for GPT-4o mini Achiam et al. (2023) across various languages and settings demonstrate that effective multilingual safeguards are attainable, contingent on the development and implementation of targeted mitigation strategies, such as comprehensive multilingual safety training datasets and more sophisticated adversarial prompting techniques for model evaluation.
@@ 5.2 Toxic QA Task
Table 3 presents the results of various LLMs’ responses when selecting an option to fill in the blank for a hateful statement in the general setting. The findings reveal sizable disparities in bias across the evaluated models. Mistral demonstrates a pronounced tendency to avoid the neutral option, consistently selecting racially targeted options across multiple low-resource languages. This behavior underscores a strong underlying bias toward racial discrimination. Similarly, SeaLLM and LLaMA exhibit biases, frequently prioritizing racially or gender-targeted options as their top two choices when the neutral option is not selected. In contrast, SEA-LION predominantly selects gender-based options over other categories when deviating from the neutral choice. Qwen and GPT-4o mini emerge as the least biased models, frequently selecting the neutral option (“I’m not sure”) as its primary response. When the neutral option is not selected, Qwen and GPT-4o mini predominantly generate “Invalid” responses, further reducing the likelihood of targeting specific social groups. These results highlight the pervasive biases embedded within LLMs when presented with all possible options, particularly in low-resource multilingual settings. The observed disparities emphasize the urgent need to develop equitable and unbiased AI systems capable of mitigating biases across diverse linguistic and cultural contexts.
@@
Our findings confirm while bias appears in English, it is notably more frequent in Malay, Tamil, and Singlish, particularly regarding race and religion-related options. Specifically, Mistral and SeaLLM consistently displayed the highest levels of bias, particularly concerning race and religion, likely attributable to insufficient cultural adaptation within their training data. SEA-LION and Llama-3.1 exhibited more inconsistent bias patterns, while Qwen and GPT-4o mini generally prioritized neutrality, albeit sometimes through evasive responses such as “Not sure”. These findings emphasize the need for continuous assessment and refinement of LLMs to address biases, particularly in low-resource and multilingual settings.
@@ 5.3 Toxic Tweet Composition Task
Our investigation confirms a strong correlation between low-resource languages and increased toxicity in LLM-generated content. Although English does not entirely escape toxic generation, the highest hateful tweet rates occur in Singlish, Malay, and Tamil. This pattern emerges across both zero-shot and few-shot learning paradigms. Specifically, Tamil exhibited among the highest toxicity rates, with models like Mistral, Qwen and Llama-3.1. Similarly, Malay demonstrated elevated toxicity, notably with Mistral and Llama-3.1. Singlish, a non-standard variety of English, further reinforces this trend, showing heightened toxicity with SeaLLM, Mistral, Qwen and Llama-3.1. These findings suggest a link between limited training data and weaker ethical alignment in low-resource languages and increased toxicity. While Chinese, a high-resource language, generally exhibited lower toxicity, the high toxicity observed with Llama-3.1 demonstrates that resource availability alone is insufficient to guarantee robust toxicity mitigation.
@@ 6 Conclusion and Future Work
Beyond academic contributions, our work has tangible real-world implications. By identifying weaknesses in LLM safety mechanisms, SGToxicGuard provides actionable insights for AI developers, policymakers, and trust & safety teams to improve content moderation systems. The dataset and the evaluation framework enable AI practitioners to improve toxicity detection in multilingual environments, thus reducing the risks of harmful AI-generated content. Furthermore, our methodology offers a scalable approach to strengthening AI safety in culturally diverse regions, ensuring that LLMs deployed in multilingual societies are both inclusive and responsible.
@@
Our research presents SGToxicGuard, a dataset and evaluation framework intended to rigorously assess Large Language Model (LLM) safety within Singapore’s low-resource, multilingual context. Purpose of the Work and Intended Use. The primary objective of this research is to identify potential failings or blind spots in LLM safety systems, not to promote or normalize hateful or toxic language. Although we use real-world examples of hateful speech, our motivation is to study how LLMs respond to and mitigate harmful content.
SGToxicGuard focuses on languages—such as Singlish, Malay, and Tamil—that may have limited resources for toxicity detection. By developing a multilingual benchmark, we aim to foster inclusivity and better protection for communities that are often overlooked in mainstream dataset creation. Our dataset includes language configurations and cultural references that reflect Singapore’s diverse population. We acknowledge nuances in local dialects, sociolects, and potential sensitivities.
This research builds on established resources (e.g., HateCheck, SGHateCheck) and does not collect new personal data from users. Identifying information has been removed or anonymized. All dataset components are released for scientific research purposes under clear licensing terms, ensuring transparency about data provenance and usage guidelines.

### v2

@@ 1 Introduction
The rapid proliferation of large language models (LLMs) has introduced unprecedented capabilities in natural language processing, enabling Artificial Intelligence (AI) applications ranging from multi-turn conversational dialogues to multilingual content generation  Dwivedi et al. (2023); Gottlieb et al. (2023); Deng et al. (2024). However, alongside these advancements, concerns regarding AI safety, ethical deployment, and content moderation have grown, particularly in managing harmful or toxic content Wang and Kim (2023); Anjum and Katarya (2024). As LLMs are increasingly integrated into real-world applications, it is paramount to ensure their responsible operation in diverse linguistic and cultural settings.
Red-teaming approaches have emerged as a prominent method for systematically probing LLM vulnerabilities Perez et al. (2022); Ganguli et al. (2022). Techniques such as adversarial input crafting Papernot et al. (2016), scenario-based evaluations Carroll (1997), and boundary condition testing Papanastasiou et al. (1992) have proven effective for detecting biases, toxic content, and other risks. However, these efforts have been largely concentrated on high-resource languages, such as English, thus leaving critical gaps in low-resource multilingual settings. Limited data availability, combined with unique linguistic variability and cultural nuances, exposes models to new vulnerabilities, increasing the risk of biased and toxic content generation in low-resource multilingual contexts. Such vulnerabilities can be exploited to perpetuate social biases and exacerbate social divides within marginalized communities. To address these challenges, there is a growing need for red-teaming methodologies that are explicitly tailored to low-resource, multilingual environments.
Singapore presents a uniquely relevant testbed for evaluating LLM safety in low-resource, multilingual contexts. The country’s linguistic landscape encompasses English, Singlish (a creole-like vernacular), Chinese, Malay, and Tamil, with frequent code-switching, culturally specific expressions, and sociolectal variations. These local linguistic factors present new challenges for AI safety evaluation that go beyond traditional toxicity detection frameworks, which were primarily designed for English. While recent efforts, such as SeaLLM Zhang et al. (2024) and SEA-LION Singapore (2024), have developed LLMs tailored to Southeast-Asian languages, the effectiveness of their safety mechanisms in minority languages remains underexplored. Existing benchmarks, such as SGHateCheck Ng et al. (2024), provide valuable functional tests for hate speech detection, but do not systematically evaluate LLM vulnerabilities under adversarial attacks. Given the increasing adoption of AI-driven productivity tools in Singapore and similar multilingual societies, developing a rigorous red-teaming evaluation framework is essential for advancing inclusive and contextually aware AI safety research.
@@
Our work makes the following key contributions: (i) We present SGToxicGuard, the first multilingual dataset designed for red-teaming LLMs in low-resource environments. (ii) We propose a three-task evaluation framework, Conversational Safety, Toxic QA, and Content Composition, that systematically assesses LLMs’ ability to handle Singapore-specific toxic content in real-world applications. (iii) Our study offers the first large-scale benchmarking of multilingual LLMs across Singapore’s low-resource languages, highlighting strengths and limitations in existing AI safety implementations.
By demonstrating the vulnerabilities of LLMs for multilingual toxicity detection, we advance AI safety for linguistically diverse societies. Our findings contribute to ongoing discussions on AI ethics, content moderation policies, and robustness in multilingual AI deployments.
@@ 2.1 Red-Teaming for AI Safety
Singapore’s linguistic landscape, which includes English, Singlish, Mandarin, Malay, and Tamil, introduces unique challenges such as code switching and cultural specificity, which existing frameworks often overlook. We address this gap by proposing a Singapore-focused red-teaming methodology to benchmark LLM safety across multilingual, real-world scenarios.
@@ 2.2 Multilingual Toxicity Detection
Recent advances in multilingual embeddings and LLMs, such as mBERT Pires et al. (2019), XLM-RoBERTa Conneau et al. (2020), GPT-3 Achiam et al. (2023), LLaMA Dubey et al. (2024), and Mistral Jiang et al. (2023), have improved cross-lingual transfer, but also amplified concerns about bias Tedeschi et al. (2024). We extend this line of research by evaluating LLM toxicity detection using SGHateCheck Ng et al. (2024), a curated dataset offering parallel toxic content across Singlish, Chinese, Malay, and Tamil. Unlike prior Southeast Asian datasets Maity et al. (2023); Lu et al. (2023); Chakravarthi et al. (2021), SGHateCheck enables structured multilingual comparisons, revealing disparities in LLM safety and informing the development of culturally sensitive AI moderation frameworks. Examples of localized hateful content containing local slurs and expressions across Singlish, Chinese, Malay, and Tamil can be found in Appendix A.6.
Beyond SGHateCheck, an expanding body of research has advanced hate speech detection across both languages and modalities, spanning academic benchmarks and real-world moderation pipelines Hee et al. (2024). For example, datasets and benchmarks such as RabakBench Chua et al. (2025) highlight the importance of localized evaluations in low-resource settings and enable cross-cultural comparisons. Recent models like LionGuard 2 Tan et al. (2025) and AngryBERT Awal et al. (2021) aim to build lightweight, emotion- and target-aware hate speech detectors, while deep representation methods like DeepHate Cao et al. (2020) and augmentation approaches such as HateGAN Cao and Lee (2020) improve robustness under data scarcity and mitigate domain shift in evolving platforms. Other studies highlight the need for adversarial robustness, particularly in dealing with cloaked and obfuscated offensive language in Chinese text Xiao et al. (2024). Additionally, model-agnostic meta-learning approaches have been proposed to facilitate multilingual adaptation while remaining mindful of practical deployment constraints Awal et al. (2023). Complementary research explores explainability and moderation workflows, including retrieval-based methods for covert toxicity detection Lee et al. (2024) and the use of LLMs to generate human-readable explanations for hateful content Hee and Lee (2025). Together, these efforts demonstrate the breadth of strategies for building robust and culturally grounded content moderation systems.
@@ 3 Framework
3.1 Toxic Conversation Task
Toxic Conversation Task evaluates LLM’s ability to generate non-toxic responses in a single-turn dialogue setting involving hateful input, which examines the model’s behavior across various social groups and low-resource languages.
This task simulates a dialogue scenario between two participants, A and B, where participant A provides a hateful statement, and participant B responds. Hateful statements from participant A are sourced from the SGHateCheck dataset Ng et al. (2024), while the LLM generates participant B’s response, which is expected to avoid toxicity by adopting constructive and respectful tones, such as refusal or dissuasion.
@@ 3.3 Toxic Tweet Composition Task
This task was evaluated under both zero-shot and two-shot in-context learning settings. In the two-shot setting, demonstration examples were provided, each showing a transformation of a hateful statement into a tweet that subtly conveys the same hateful message. These examples, while crafted to simulate malicious intent, are used to assess the LLM’s vulnerability to generating subtle hateful content. Additionally, they help evaluate whether exposure to such examples could trigger a “toxicity jailbreak”, where the model deviates from its intended ethical safeguards. We also examine cross-lingual transfer by translating demonstrations and probing whether stylistic cues generalize across languages, revealing gaps between safety alignment and compositional generalization. Examples of tweet composition and demonstration are listed in Appendix A.2 and A.3.
@@ 5.1 Toxic Conversation Task
We can see in Table 2 that, in the zero-shot setting, most LLMs demonstrate robust safety mechanisms, producing relatively few hateful responses across different languages and the models exhibit lowest hateful response rate in English. However, notable exceptions include Mistral Jiang et al. (2023) and SeaLLM Zhang et al. (2024), which exhibit vulnerabilities in specific languages. Mistral generates a sizable proportion of hateful responses in Malay (9.74%) and Tamil (7.86%), underscoring potential weaknesses in handling low-resource languages. Surprisingly, the specialized multilingual LLM, SeaLLM, performs worse, with higher rates of hateful responses in Singlish (8.82%), Chinese (7.46%), and Tamil (14.64%). These results indicate that models fine-tuned for Southeast Asian languages face considerable challenges in mitigating toxicity in low-resource contexts.
@@
Introducing two hateful dialogue demonstrations in the two-shot setting results in a notable increase in the generation of hateful responses across all LLMs and languages. Qwen Team (2024) demonstrates the highest resilience to in-context learning, maintaining the lowest overall rate of hateful responses. However, even Qwen exhibits a sizable increase in Tamil (22.72%). In contrast, Mistral Jiang et al. (2023) and Llama-3.1 Dubey et al. (2024) show moderate increase in hateful response rates across languages, ranging from approximately 2% to 14%.
SeaLLM, despite higher rates in the zero-shot setting, demonstrates relatively lower rates in Malay (2.66%) and Chinese (11.01%) under two-shot conditions. SEA-LION Singapore (2024), however, exhibits alarming susceptibility to toxicity induction, with hateful response rates increasing to 61.52% in Chinese and 58.42% in Tamil. These results suggest that in-context learning substantially amplifies the risk of toxicity jailbreaks in certain models.
@@
Our result reveals a major increase in toxic content generation when interacting with users of low-resource languages, under adversarial prompting conditions. In zero-shot evaluations, we observed elevated toxicity levels in Malay, Tamil, and Chinese, suggesting limitations in the safety alignment of multilingual LLMs. This trend was amplified in two-shot scenarios, with substantial increase in toxicity for Tamil in SEA-LION Singapore (2024) and Qwen Team (2024). Such differences suggest that existing safeguards may be more thoroughly tuned for English data. However, the low toxicity rates observed for GPT-4o mini Achiam et al. (2023) across various languages demonstrate that effective multilingual safeguards are attainable, contingent on the implementation of targeted mitigation strategies, such as comprehensive multilingual safety training datasets and more complicated adversarial prompting techniques for model evaluation.
@@ 5.2 Toxic QA Task
Table 3 presents results of LLMs selecting options to fill blanks for hateful statements in the general setting. The findings reveal disparities in bias across the evaluated models. Mistral tends to avoid the neutral option, consistently selecting racially targeted options across multiple low-resource languages, underscoring a strong bias toward racial discrimination. Similarly, SeaLLM Zhang et al. (2024) and LLaMA Dubey et al. (2024) prioritize racially or gender-targeted options as their top two choices when the neutral option is not selected.
In contrast, SEA-LION more often selects gender-based options when deviating from neutral. Qwen and GPT-4o mini are the least biased, frequently selecting the neutral option (‘I’m not sure”) as the primary response. When the neutral option is not selected, they predominantly generate ‘Invalid” responses, further reducing the likelihood of targeting specific social groups.
These results highlight pervasive biases in LLMs when presented with all options, particularly in low-resource multilingual settings. The disparities emphasize the urgent need to develop equitable, unbiased AI systems capable of mitigating biases across diverse linguistic and cultural contexts.
@@
Our findings confirm bias appears in English but is higher in Malay, Tamil, and Singlish, especially in race and religion related options. Mistral and SeaLLM showed the highest bias on race and religion, likely due to limited cultural adaptation in training data. SEA-LION and Llama-3.1 showed inconsistent patterns, while Qwen and GPT-4o mini prioritized neutrality, sometimes via evasive responses like “Not sure”. These findings underscore the need for continuous assessment and refinement of LLMs to address biases, especially in low-resource and multilingual settings.
@@ 5.3 Toxic Tweet Composition Task
Our investigation confirms a strong correlation between low-resource languages and increased toxicity in LLM-generated content. This effect remains pronounced across model families and parameter scales. Although English does not entirely escape toxic generation, the highest hateful tweet rates occur in Singlish, Malay, and Tamil. This pattern emerges across both zero-shot and few-shot learning paradigms, and persists under varied prompts and topics. Specifically, Tamil exhibited among the highest toxicity rates, with models like Mistral, Qwen and Llama-3.1. Similarly, Malay demonstrated elevated toxicity, notably with Mistral and Llama-3.1. Singlish, a non-standard variety of English, further reinforces this trend, showing heightened toxicity with SeaLLM, Mistral, Qwen and Llama-3.1.
These findings suggest a link between limited training data and weaker ethical alignment in low-resource languages and increased toxicity. While Chinese, a high-resource language, generally exhibited lower toxicity, the high toxicity observed with Llama-3.1 demonstrates that resource availability alone is insufficient to guarantee robust toxicity mitigation.
@@ 6 Conclusion and Future Work
Beyond academia, our work has real-world impact. By identifying weaknesses in LLM safety mechanisms, SGToxicGuard provides insights for AI developers, policymakers, and safety teams to strengthen moderation. The dataset and evaluation framework help AI practitioners improve toxicity detection in multilingual settings, reducing risks of harmful AI-generated content. Our methodology offers a scalable way to strengthen AI safety in culturally diverse regions, ensuring LLMs deployed in multilingual societies are inclusive and responsible.
@@
Our research presents SGToxicGuard, a dataset and evaluation framework intended to assess Large Language Model (LLM) safety within Singapore’s low-resource, multilingual context. Purpose of the Work and Intended Use.
The primary objective of this research is to identify potential failings or blind spots in LLM safety systems, not to promote or normalize hateful or toxic language. Although we use real-world examples of hateful speech, our motivation is to study how LLMs respond to and mitigate harmful content.
SGToxicGuard focuses on languages, such as Singlish, Malay, and Tamil, that may have limited resources for toxicity detection. By developing a multilingual benchmark, we aim to foster inclusivity and better protection for communities that are often overlooked in mainstream dataset creation. Our dataset includes language configurations and cultural references that reflect Singapore’s diverse population. We acknowledge nuances in local dialects, sociolects, and potential sensitivities.
This research builds on established resources and does not collect new personal data from users. Identifying information has been removed or anonymized. All dataset components are released for scientific research purposes under clear licensing terms, ensuring transparency about data provenance and usage guidelines.